OpenStudy (anonymous):

Write the equation in standard form: y=-1/4x+6

OpenStudy (helder_edwin):

multiply what you have by -4 what do you get?

OpenStudy (helder_edwin):

no wait just put 6 to the other side of the equation

OpenStudy (anonymous):

x+4y=24

OpenStudy (helder_edwin):

you get \[ y-6=-\frac{1}{4}x \] which has the form \[ y-k=m(x-h) \]